Sol Summary – February 22nd

Crew 293 Sol Summary Report 22-02-2024

Sol: 4

Summary Title: First clouds upon the Crew

Author’s name: Marie Delaroche, Crew Commander

Mission Status: Nominal

Sol Activity Summary:

This morning, the second EVA team tasked with installing the remaining atmospheric devices (the Field Mill and a weather station) left the station at 9:15. The team successfully installed a weather station which is now synchronized with a tablet inside the Hab. The data collected by the weather station will be used to determine how the measurements taken by the atmospheric instruments are influenced by wind speed, wind direction, humidity, etc.The rest of the EVA went smoothly, apart from a malfunction with the Field Mill.

One EVA crewmember started noticing an issue with the airflow of their suit as the EVA crew was approaching the airlock. Once inside, they felt their suit was no longer ventilated, and quickly asked the crew to take off their helmet, which they did immediately; sim was broken for safety purposes. The crewmember sat down and had difficulty breathing, as they had been breathing accumulated CO2. The entire crew reacted to the situation perfectly; we called Mission Support, and provided all the correct first aid elements. Mathurin checked their oximetry level, which was back to normal after they went back to the Hab. Once the situation was under control, Leo and Marie performed measurements on the ventilation battery and noticed that the battery level had dropped below 8V, after 3h30mins of use.

After re-reading previous emails between Mission Support and Crew 292, we realized the faulty suit (suit 6) was also reported to have battery issues, and that the battery had already been replaced before. We will make sure that the problem does not occur again by testing every suit (see Operations Report for details on the tests) to identify the ones that might be faulty.

The Crew spent the early afternoon resting after an eventful morning; we then spent time training with the AMI interface, and troubleshooting sensors.

Look Ahead Plan:

Tomorrow will be dedicated to various activities, most notably preparing videos for the schools linked to our outreach program, continuing the EchoFinder experiment, and troubleshooting hardware for Orbital Architecture.

An EVA to retrieve and fix the malfunctioning Field Mill is scheduled on Sol 6.

Anomalies in work: EVA Suit #6

Weather: Sunny, High 15°C, Low 5°C

Crew Physical Status: One crewmember had trouble breathing at the end of the EVA. They have fully recovered.

EVA: Deployment of two Atmospheric Instruments, the Field Mill and a weather station.

Reports to be filed: Sol Summary, Journalist’s Report, GreenHab Report, Operations Report, Photos (6-8 pics), EVA Reports (1), HSO Report

Support Requested: NA

Copyright © The Mars Society. All rights reserved. | Main Site